OpenStudy (anonymous):

anyone help me Write the explicit formula for the geometric sequence. a1 = -5 a2 = 20 a3 = -80

Nnesha (nnesha):

what's the formula for geometric sequence ??

OpenStudy (anonymous):

For r\neq 1, the sum of the first n terms of a geometric series is: a + ar + a r^2 + a r^3 + \cdots + a r^{n-1} = \sum_{k=0}^{n-1} ar^k= a \, \frac{1-r^{n}}{1-r},

Nnesha (nnesha):

\[a_n = a_1 \times (r)^{n-1}\] formula where a_1 is first term and r is ratio

Nnesha (nnesha):

what's the ratio for that sequence ? to find ratio \[\huge\rm r = \frac{ a_2 }{ a_1 }\] divide next term by previous one 2nd term/1st term

OpenStudy (mathteacher1729):

If it's asking for an explicit formula, it probably wants you to identify \(a\) and \(r\) then use summation notation instead of writing the series as a sum of individual terms: \(\sum_{n=1}^{3}a_n\) where \(a_n=a\cdot r^n\).

OpenStudy (anonymous):

okay

Nnesha (nnesha):

i guess they just need formula a_1 is already there

OpenStudy (anonymous):

thax for the help

Nnesha (nnesha):

what's "r" equal to ?? r= ??

OpenStudy (anonymous):

i will find out i know all this just that i was a liitle cunfused

Nnesha (nnesha):

okay... let me know what u get :-) take ur time

OpenStudy (anonymous):

here are the answers A an = -5* (-4)^n B an = -5(-4)^n-1 C an = -4(-5)^n-1 i got D wrong

Nnesha (nnesha):

that isn't answer of my questions :(

Nnesha (nnesha):

okay so r = ???

OpenStudy (mathteacher1729):

Ok, back. So we have alternating positive and negative numbers, so you know you have a \((-1)^n\) in the explicit formula. Now 5, 20, 80... all have a factor of 5 in common: 5(1) =5 5(4) =20 5(16) = 80 Does this help?